Creativeone24:

Creativeone24 Jason paid $15.50 for 3 slices of pizza and 2 burgers. Susan paid $20 for 1 slice of pizza and 4 burgers. Fill in the following:(Do NOT include a $ in your answer) Each slice of pizza costs $ _______________________________ . Each burger costs $ _______________________

Creativeone24:

@dude

Creativeone24:

Creativeone24:

Vocaloid:

I'm really sorry but if you're waiting for someone to do the problem for you, it's against the rules to give direct answers here I have tried to help you several times, but I can't do the work for you :/

Creativeone24:

😭😭😭😭 sorry I thought you was gone I’m new here

Creativeone24:

Vacaloid

Creativeone24:

@Vacaloid

Creativeone24:

I also decided lI had to post the question up again

Vocaloid:

let's start from the beginning again: 3p + 2b = 15.50 Susan's order: p + 4b = 20 solve for p and b using elimination or substitution start by multiplying equation 1 by 2 let me know what you get

Vocaloid:

if you are having trouble, that is perfectly fine, but I can't read your mind, you have to say something.

Creativeone24:

ok im working on it

Creativeone24:

@Valacoid p=6.2

Vocaloid:

that is not correct what did you get after multiplying 3p + 2b = 15.50 by 2?

Vocaloid:

just take every term in this equation and multiply it by 2 3p * 2 = ?

Creativeone24:

im almost finish i know its a long hold up

Creativeone24:

6p+4p=31 is ths correct

Vocaloid:

good (should be a 4b term not a p term though) so 6p + 4b = 31 p + 4b = 20 now subtract the two equations 6p - p = ?

Creativeone24:

tht will = 5p?

Creativeone24:

im correct? or do i have to do both sides?

Vocaloid:

good, and 31 - 20 = 11 now, 5p = 11 p = ?

Creativeone24:

ok

Vocaloid:

5p = 11 divide both sides by 5 to solve for p.

Creativeone24:

p=2.2

Vocaloid:

awesome, now let's plug this back into equation 2 and solve for b p + 4b = 20 so 2.2 + 4b = 20 what does b equal?

Creativeone24:

b=4.45 ??

Vocaloid:

awesome! so p = 2.2 and b = 4.45 and those are your answers

Creativeone24:

thnk u very much very much i appreciate it !!!

Creativeone24:

@Vocaloid

Creativeone24:

Tomorrow if i need help ill let you know gb